OpenStudy (bbb911):

This is GEometry: Solve the following system of equations by graphing y=x-4 2y=-x+10

OpenStudy (anonymous):

I LOVE THESE! I'll give you the algebraic answer.

OpenStudy (bbb911):

yayyyyyyyy :D

OpenStudy (anonymous):

You can graph it by doing this: y=x-4 y = (-x +10)/2 or: y = -1/2x + 5 I'll explain it via algebra next. :P

OpenStudy (anonymous):

since you know y = x-4, plug the (x-4) as your Y value in the second equation: 2(x-4) = -x + 10 2x - 8 = -x + 10 2x = -x + 18 3x = 18 x = 6 BOO YA! (lemme double check this first. :P)

OpenStudy (bbb911):

oh ok

OpenStudy (anonymous):

So, the point of intersection should be (6,2). The two came from plugging in 6 and doing 6-4=y and y = 2. We test from here: 2y=-x+10 2(2) = -(6)+10 4 = 4 It's true. There's your answer. (6,2) ^_^

OpenStudy (bbb911):

Thank you so much! :3

OpenStudy (anonymous):

No problem, ma'am! Thanks for the trophy. Keep the trophies comin' and I'll keep the answers coming. :P

OpenStudy (bbb911):

yay!!!!!!!!!!! i like the way you think